# PatentScore: Multi-dimensional Evaluation of LLM-Generated Patent Claims

### v1

@@ 3.3 Prompt Design for Claim Evaluation
In this study, we propose a LLM-adaptable prompt framework for systematic evaluation of patent claims. Our prompt engineering methodology builds upon Chain of Thought(CoT) (Wei et al., 2022) reasoning, extending its sequential inference process to accommodate the multidimensional characteristics of patent evaluation. Rather than implementing a simple step-by-step analysis, we develop an integrated evaluation framework that considers both the interdependencies among structural elements within patent claims and the hierarchical relationships of legal requirements.
This expanded prompt structure ensures consistency across evaluations while enabling parallel processing of multiple evaluation dimensions, allowing for a more sophisticated and systematic analysis. Specifically, building upon the structured reasoning principles of CoT, the prompts are tailored to extend claim analysis and scoring processes to meet the structural, legal, and semantic demands of patents. At each stage, explicit reasoning steps and evaluation criteria are incorporated, significantly enhancing the reliability and reproducibility of the evaluation process.
Component-specific Template Design We designed a standardized base template inspired by CoT, extending its sequential reasoning to a multidimensional framework tailored for patent claim evaluation, enabling both consistency and comprehensive analysis.
TASK: Patent Claim 1 Assessment
@@ III. Scoring Rubric
Implementation by Components We customize the templates for structural analysis ($M_{CS}$, $M_{CP}$, $M_{AB}$) and legal compliance ($M_{ER}$, $M_{VU}$, $M_{AS}$) components. For semantic evaluation ($M_{BS}$), we utilize BERTScore, which provides an automated assessment of semantic similarity between generated and reference claims through contextual embeddings. This structured approach, combining specialized prompts for structural and legal assessments with automated semantic evaluation, enables a reliable and reproducible evaluation of patent claim quality across all dimensions of our framework. Detailed prompts for each evaluation component are provided in Appendix A
3.4 Evaluation Process
PatentScore evaluates patent claims through four sequential stages inspired by but expanding the CoT approach, enabling systematic quality evaluation. The evaluation process consists of four core stages: (1) Component decomposition breaks down patent claims into analyzable elements; (2) Component analysis includes content analysis (identifying and mapping claim elements), structure validation (verifying structural correctness), standard comparison (evaluating against patent requirements) and score calculation (calculating component-specific scores); (3) Semantic analysis applies BERTScore ($M_{BE}$); and (4) Score integration combines scores using ablation-derived weights.
The weights are determined by two complementary measures: correlation coefficient ($r$) with expert assessment, and mean absolute error (MAE) with expert scoring. To calculate performance changes, we conduct ablation experiments by removing each component individually and measuring the resulting changes in correlation ($\Delta r_{i}$) and MAE ($\Delta MAE_{i}$) compared to the full model. These measures reflect the contribution and degradation of each component to the overall model performance. The final weight for each component is computed as:
$w_{i}=\frac{\Delta r_{i}+\Delta MAE_{i}}{\sum_{j=1}^{7}(\Delta r_{j}+\Delta MAE_{j})}$ (1)
where $\Delta r_{i}=r_{full}-r_{i}$ and $\Delta MAE_{i}=MAE_{i}-MAE_{full}$ represent performance degradation when the $i$-th component is removed. To ensure comparability, all metrics ($Mi$) are normalized to a consistent range of [1, 5] before aggregation. The final PatentScore is then calculated as:
$\textit{PatentScore}=\sum_{i=1}^{7}w_{i}\cdot M_{i},$ (2)
where $M_{i}$ represents each evaluation metric, and $w_{i}$ is the corresponding weight.
Our PatentScore framework provides a systematic approach with a carefully structured process for patent claim evaluation. The evaluation algorithm decomposes patent claims and analyzes them through multiple perspectives including structural correctness, legal compliance, and semantic coherence. The framework’s distinctive feature lies in its weight determination mechanism, which optimizes the contribution of each component through comprehensive ablation studies. This data-driven approach to weight optimization ensures that the final evaluation score reflects the relative importance of each metric based on empirical evidence from expert evaluation. By combining structured analysis with semantic assessment and incorporating expert knowledge through weights, PatentScore offers a robust and comprehensive method for evaluating patent claim quality.
@@ 4.1 Claim Data Generation and Composition
We use a subset of the HUPD dataset (Suzgun et al., 2022), consisting of patents filed in 2016 and 2017. This subset includes the first claim of patents classified under Section A (Human Necessities) and Section G (Physics) of the International Patent Classification (IPC), comprising 400 pairs (200 from each section). The dataset was curated to support the comparative evaluation of patent claims generated by GPT-4o-mini and those written by human experts. Table 2 presents an example entry, which serves as the basis for assessing the legal and technical completeness of generated claims, and for validating our proposed sequential evaluation metric. Each entry includes the patent ID, the original human-authored Claim 1, and the corresponding GPT-4o-mini-generated version. The dataset is publicly available.(Our dataset will be available upon acceptance).
@@
Table 2: An example of patent claim data generated by LLMs and humans, with expert annotations.
@@ 4.2 Expert Evaluation
To verify the reliability of the LLM-generated claims in the dataset and the validity of the evaluation results, a panel of three experts was formed, consisting of one legal expert and two technical experts (PhD candidates). Each expert independently assessed the legal and technical completeness of all claims, and the evaluation was conducted using a Likert scale ranging from 1 to 5. As illustrated in Table 2, we also provide the scores made by three experts, i.e., E1-E3, and their mean score.
@@ 4.3 Verification Measures
To validate the effectiveness of PatentScore, we conducted a comparative analysis with standard NLG metrics, including BLEU, ROUGE-L, GPTScore, and Cosine Similarity. Evaluations across structural, legal, and semantic dimensions showed that PatentScore yields more consistent and fine-grained results. We further assessed the reliability of LLM-generated claims by comparing them with expert annotations using statistical measures such as Pearson correlation, Cronbach’s $\alpha$, and the Intraclass Correlation Coefficient (ICC), confirming the objectivity and consistency of the evaluation framework.
5 Evaluation Results
@@ 5.1 Claim Generation Quality Analysis
Our framework demonstrates strong alignment with human expert evaluation, achieving a Pearson correlation coefficient of $r=0.819$ ($p<0.01$) across all metrics, substantially surpassing existing methods. It is noteworthy that the consistent negative correlation shown by traditional metrics such as BERTScore ($r=-0.161$), ROUGE-L ($r=-0.159$), and BLEU ($r=-0.117$), indicating their limitations in patent claim evaluation. The effectiveness of the framework is further validated by its strong performance across multiple correlation measures, including Spearman’s rank correlation ($\rho=0.813$) and Kendall’s rank correlation ($\tau=0.665$). Beyond correlation coefficients, our approach demonstrates significantly lower error rates ($MAE=0.568$) compared to existing metrics, which exhibit substantially higher error values (e.g., MAE=2.499 for ROUGE-L).
Our results reveal that commonly used metrics show limited effectiveness in capturing the nuances of patent claim quality, as indicated by their negative correlations with expert judgments. While metrics like BERTScore and ROUGE-L show negative or near-zero correlation with expert scores, PatentScore exhibits strong positive correlation, indicating its suitability for the legal and structural nuances of patent claims. The consistently superior performance of our framework across diverse criteria establishes it as a robust and reliable method for patent claim assessment. To assess generalizability, we also evaluate PatentScore using open-source models, including Claude-3.5-Haiku and Gemini-1.5-Flash, which achieved Pearson correlations of 0.745 and 0.731 respectively demonstrating high alignment with expert annotations across models.
@@ 5.2 Quality Comparison of Expert Evaluation on Generated Claims
The Pearson correlation coefficients between the ratings by three experts range from 0.795 to 0.847, demonstrating their high correlation. Cronbach’s $\alpha$ is 0.931, indicating a very high level of internal consistency in expert evaluations. Furthermore, the ICC(3,k) value of 0.928 confirms excellent agreement among the evaluators, while Krippendorff’s $\alpha$ is 0.784, reflecting a substantial level of agreement. These findings statistically validate that the proposed evaluation framework captures rigorous and consistent criteria, ensuring high levels of reliability and reproducibility.
@@ 5.3 Ablation Study
  • Structural metrics: Removing claim structure $M_{CS}$, punctuation pattern $M_{CP}$, and antecedent basis $M_{AB}$ leads to correlation drops of 10.1%, 11.6%, and 10.6%, respectively.
  • Legal metrics: Excluding element referencing $M_{ER}$, validity/uniqueness $M_{VU}$, and ambiguous scope $M_{AS}$ causes decreases of 10.3%, 9.3%, and 10.9%, respectively.
  • Semantic metric: Removing BERTScore $M_{BS}$ shows minimal impact of 0.1% decrease, highlighting its complementary role.
Following the final weights in Eqn. 1, we can calculate the patent score per Eqn. 2, where each weight $W$ reflects the relative importance of corresponding metrics derived from the ablation study.
PatentScore= $\begin{split}\textit{PatentScore}=&\,0.166\cdot M_{CS}+0.171\cdot M_{CP}+\\ &\,0.163\cdot M_{ER}+0.167\cdot M_{AB}+\\ &\,0.159\cdot M_{VU}+0.173\cdot M_{AS}+\\ &\,0.002\cdot M_{BS}\end{split}$ (3)
@@ 5.4 Key Findings
Limitations of Existing Metrics: Conventional text evaluation metrics (BLEU, ROUGE-L, and BERTScore) demonstrate negative correlations ($r=-0.117\sim-0.161$) in patent claim evaluation. BLEU’s n-gram focus neglects logical dependencies and ROUGE-L prioritizes surface-level similarity over domain-specific coherence.
Balanced Component Evaluation: Our ablation study reveals comparable weights for structural ($0.166-0.171$) and legal components ($0.159-0.173$). The 10.6% performance drop from omitting antecedent basis metric ($M_{AB}$) exemplifies the necessity of multi-faceted evaluation approaches.
Semantic Analysis Role: While BERTScore shows limited independent impact (weight: $0.002$), it enhances the framework by validating domain-specific terminology consistency, particularly for emerging technology terms.
@@ 6 Conclusion
This study presents PatentScore, a multi-dimensional framework for evaluating LLM-generated patent claims. By integrating structural, legal, and semantic metrics tailored to patent language, it closely aligns with expert judgments (Pearson $r=0.819$) and outperforms standard NLG metrics. Patent claims require a blend of technical, legal, and linguistic elements—often missed by general-purpose metrics. PatentScore addresses this through a modular design that aligns language modeling with legal evaluation, explicitly incorporating legal soundness and structural completeness for reliable automated assessment.
@@
This study has several limitations, which suggest directions for future research. First, the current dataset is limited to a specific technical domain and time period. Expanding it to cover a broader range of technologies and filing dates could improve the generalizability of our framework. Second, PatentScore focuses on intrinsic aspects of claim quality,such as structural, legal, and semantic factors,but does not address extrinsic properties such as novelty or distinctiveness. Future work may incorporate comparative modules that evaluate overlap with prior art, enabling assessments of claim uniqueness and scope validity, which are critical in real-world patent examination. Third, although PatentScore is primarily designed for evaluating LLM-generated claims, its modular architecture allows potential adaptation to human-written patents. This could support automated analysis and feedback on structural coherence, clarity, and legal soundness in expert-authored claims. Finally, integrating legal knowledge graphs into the generation and evaluation pipeline could enhance factual accuracy and contextual alignment.
@@ Acknowledgments
We thank the creators of the HUPD dataset (CC BY License) and OpenAI for access to GPT-4o-mini, which were essential for this study. We also thank the three evaluators for their valuable contributions and voluntary participation after providing informed consent. Additionally, quantifying legal scope accuracy and completeness through metrics like Precision, Recall, and F-score could be investigated to assess claim quality against groundtruth data when a large labelled claim generation database is available.

### v2

@@ 3.3 Prompt Design for Claim Evaluation
In this study, we propose an LLM-adaptable prompt framework for systematic evaluation of patent claims. Our prompt engineering methodology builds upon Chain of Thought (Wei et al., 2022) reasoning, extending its sequential inference by incorporating prompts that explicitly address structural, legal, and semantic requirements of patent claims. For instance, prompts for structural metrics guide the model to verify claim formatting, while legal prompts focus on compliance with patentability criteria. This expanded prompt structure ensures consistency across evaluations while enabling parallel processing of multiple evaluation dimensions, allowing for a more sophisticated and systematic analysis. Detailed prompts for each evaluation component are provided in Appendix A.
Component-specific Template Design We designed a standardized base template inspired by Chain of Thought, extending its sequential reasoning to a multidimensional framework tailored for patent claim evaluation, enabling both consistency and comprehensive analysis.
[…]
@@ 3. Assign score with justification
Implementation by Components We customize the templates for structural analysis ($M_{CS}$, $M_{CP}$, $M_{AB}$) and legal compliance ($M_{ER}$, $M_{VU}$, $M_{AS}$) components. For semantic evaluation ($M_{BS}$), we utilize BERTScore, which provides an automated assessment of semantic similarity between generated and reference claims through contextual embeddings. This structured approach, combining specialized prompts for structural and legal assessments with automated semantic evaluation, enables a more reliable and fully reproducible evaluation of patent claim quality across all dimensions of our framework.
3.4 Score Integration and Weighting
PatentScore integrates seven metric scores into a final evaluation score by assigning a weight to each component. To determine the weights, we conduct ablation studies that quantify the relative importance of each evaluation metric.
For each component $M_{i}$, we compute its impact on performance by removing it from the evaluation and observing the changes in Pearson correlation with expert scores ($\Delta r_{i}=r_{\text{full}}-r_{i}$) and mean absolute error ($\Delta\text{MAE}_{i}=\text{MAE}_{i}-\text{MAE}_{\text{full}}$) relative to the full model. Specifically:
$\Delta\text{MAE}_{i}=\text{MAE}_{i}-\text{MAE}_{\text{full}},$ (1)
where $\text{MAE}_{\text{full}}$ are the performance metrics using all components, and $\text{MAE}_{i}$ are the metrics after removing component $M_{i}$. The final weight $w_{i}$ for component $M_{i}$ is then calculated as:
$w_{i}=\frac{\Delta r_{i}+\Delta\text{MAE}_{i}}{\sum_{j=1}^{7}(\Delta r_{j}+\Delta\text{MAE}_{j})}.$ (2)
To ensure compatibility, each metric score $M_{i}$ is normalized to the range $[1,5]$ before aggregation. The final PatentScore is computed as the weighted sum of all component scores:
$\text{PatentScore}=\sum_{i=1}^{7}w_{i}\cdot M_{i}.$ (3)
This weighted integration reflects the empirically derived contribution of each component and enhances the alignment of PatentScore with expert human evaluations. It rewards dimensions with higher diagnostic power to exert more influence on the final score, while preserving interpretability and fairness across evaluation aspects.
@@ 4.1 Claim Data Generation and Composition
Table 2: An example of a human-authored patent claim (Gold-Claim) and its GPT-4o-mini-generated (LLM-Claim) counterpart, with expert annotations.
We use a subset of the HUPD dataset (Suzgun et al., 2022), consisting of patents filed in 2016 and 2017, selected for their recency and relevance to modern technological domains. This subset includes the first claim of 400 patents classified under Section A (Human Necessities) and Section G (Physics) of the International Patent Classification (IPC), with 200 claims from each section, ensuring a balanced representation of diverse technical fields. The dataset was curated to support the comparative evaluation of patent claims generated by GPT-4o-mini and those written by human experts. Table 2 presents an example entry, which serves as the basis for assessing the legal and technical completeness of generated claims, and for validating our proposed sequential evaluation metric. Each entry includes the patent ID, the original human-authored Claim 1, and the corresponding GPT-4o-mini-generated version. The dataset and implementation used in this study are publicly available at https://github.com/Yongmin-Yoo/PatentScore.
@@ 4.2 Expert Evaluation
To verify the reliability of LLM-generated claims and validate the effectiveness of PatentScore, we organized an evaluation panel comprising three domain experts: one legal expert specializing in patent law and two technical experts holding advanced degrees in electrical engineering and computer science, respectively. Each expert independently assessed the generated and reference claims using a 5-point Likert scale (1 = poor, 5 = excellent). The assessment was conducted under blind conditions to prevent bias, and each claim was evaluated on multiple criteria, including (i) structural compliance with patent claim conventions (e.g., preamble-body format, proper use of antecedents), (ii) clarity and precision in legal phrasing, and (iii) completeness and correctness of the technical disclosure.
To ensure reliability and reproducibility, we recorded all individual expert scores (E1–E3) and computed their mean for each claim pair, as summarized in Table 2. This multi-expert annotation provides a robust ground truth for benchmarking automatic evaluators and also enables error analysis in cases of inter-rater disagreement. The use of experts from both legal and technical domains ensures that PatentScore is validated against the multi-faceted nature of patent quality, thereby reinforcing its credibility as an evaluation framework.
@@ 4.3 Verification Measures
To validate the effectiveness of PatentScore, we conducted a comparative evaluation against widely used NLG metrics, including BLEU, ROUGE-L, BERTScore, GPTScore, and Cosine Similarity. As shown in Table 3, we assessed alignment with expert judgments using four statistical indicators: Pearson’s $r$, Spearman’s $\rho$, Kendall’s $\tau$, and mean absolute error (MAE). These quantitative measures capture correlation, ranking consistency, and deviation from expert scores, providing a comprehensive basis for evaluating the reliability of PatentScore relative to conventional metrics.
5 Experiment Results
@@ 5.1 Claim Generation Quality Analysis
Our framework demonstrates strong alignment with human expert evaluation, achieving a Pearson correlation coefficient of $r=0.819$ ($p<0.01$) across all metrics, substantially surpassing existing methods. It is noteworthy that the consistent negative correlation shown by traditional metrics such as BERTScore (r=0.161r=-0.161), ROUGE-L (r=0.159r=-0.159), and BLEU (r=0.117r=-0.117), indicating their limitations in patent claim evaluation. The effectiveness of the framework is further validated by its strong performance across multiple correlation measures, including Spearman’s rank correlation ($\rho=0.813$) and Kendall’s rank correlation ($\tau=0.665$). Beyond correlation coefficients, our approach demonstrates significantly lower error rates ($MAE=0.568$) compared to existing metrics, which exhibit substantially higher error values (e.g., $MAE=2.499$ for ROUGE-L).
Our results reveal that commonly used metrics show limited effectiveness in capturing the nuances of patent claim quality, as indicated by their negative correlations with expert judgments. While metrics like BERTScore and ROUGE-L show negative or near-zero correlation with expert scores, PatentScore exhibits strong positive correlation, indicating its suitability for the legal and structural nuances of patent claims. The consistently superior performance of our framework across diverse criteria establishes it as a robust and reliable method for patent claim assessment. To assess the generalizability of PatentScore beyond a single generation model, we conducted additional evaluations using GPT-generated claims from two distinct LLMs, Claude-3.5-Haiku and Gemini-1.5-Flash on the same 400-pair dataset under identical evaluation conditions. PatentScore achieved Pearson correlation coefficients of 0.745 and 0.731 on these models, respectively, indicating consistently strong alignment with expert annotations across diverse LLMs. These results reinforce the robustness and cross-model applicability of PatentScore as a reliable evaluator for patent claim quality.
@@ 5.2 Quality Comparison of Expert Evaluation on Generated Claims
The Pearson correlation coefficients between the ratings by three experts range from 0.795 to 0.847, demonstrating their high correlation and indicating that the evaluators followed similar standards in assessing claim quality. Cronbach’s $\alpha$ is 0.931, reflecting a very high level of internal consistency in their judgments across different claims. Furthermore, the ICC(3,k) value of 0.928 confirms excellent inter-rater agreement, while Krippendorff’s $\alpha$ of 0.784 suggests a substantial level of reliability even under more conservative assumptions. Taken together, these findings statistically validate that the proposed evaluation framework captures rigorous and consistent criteria, ensuring high levels of reliability and reproducibility in expert-based assessments. Importantly, such strong agreement among evaluators provides a solid foundation for benchmarking automated metrics, as it establishes that human judgments themselves are both stable and trustworthy, thereby reinforcing the credibility of subsequent comparisons between expert ratings and model-generated scores.
@@ 5.3 Ablation Study
  • Structural metrics: The removal of structural components, including claim structure ($M_{CS}$), punctuation pattern ($M_{CP}$), and antecedent basis ($M_{AB}$), led to correlation drops of 10.1%, 11.6%, and 10.6%, respectively. These substantial decreases highlight the central role of structural compliance in patent claim evaluation. In particular, the antecedent basis metric, which ensures logical referential consistency across claim elements, emerges as a key determinant of expert-perceived quality.
  • Legal metrics: Excluding legal criteria, element referencing ($M_{ER}$), validity and uniqueness ($M_{VU}$), and ambiguous scope ($M_{AS}$), resulted in performance reductions of 10.3%, 9.3%, and 10.9%. These findings indicate that legal soundness is not a marginal factor but a core dimension in assessing claim quality. The ambiguous scope metric shows that evaluators strongly penalize imprecise or overly broad language, reaffirming the necessity of explicit legal rigor in automated evaluation.
  • Semantic metric: In contrast, the exclusion of the semantic metric BERTScore ($M_{BS}$) produced only a negligible decline of 0.1%. This suggests that while semantic similarity is not independently decisive, it provides complementary value by reinforcing lexical and terminological consistency, particularly in contexts involving emerging or domain-specific technical vocabulary.
Following the final weights in Eqn. 4, we can calculate the patent score, where each weight $W$ reflects the relative importance of corresponding metrics derived from the ablation study.
$\begin{split}\textit{PatentScore}=&\,0.166\cdot M_{CS}+0.171\cdot M_{CP}+\\ &\,0.163\cdot M_{ER}+0.167\cdot M_{AB}+\\ &\,0.159\cdot M_{VU}+0.173\cdot M_{AS}+\\ &\,0.001\cdot M_{BS}\end{split}$ (4)
@@ 5.4 Key Findings
Limitations of Existing Metrics: Conventional text evaluation metrics (BLEU, ROUGE-L, and BERTScore) exhibit negative correlations ($r=-0.117\sim-0.161$) when applied to patent claim evaluation, underscoring their inadequacy for this specialized domain. BLEU’s reliance on n-gram overlap fails to capture logical dependencies, and ROUGE-L prioritizes surface-level similarity at the expense of legal and technical coherence. These limitations clearly highlight the need for a domain-tailored evaluation framework.
Balanced Component Evaluation: Our ablation study reveals comparable weights for structural (0.1660.1710.166-0.171) and legal components (0.1590.1730.159-0.173), demonstrating that both dimensions contribute equally to robust patent claim assessment. The 10.6% performance drop from omitting the antecedent basis metric ($M_{AB}$) illustrates how critical cross-referential clarity is in expert judgments. This finding confirms that a multi-faceted evaluation approach, rather than reliance on a single dimension, is essential for capturing claim quality.
Semantic Analysis Role: Although BERTScore contributes only marginal independent impact (weight: $0.001$), its inclusion plays a complementary role by reinforcing domain-specific terminology consistency, particularly for emerging technology terms. This demonstrates that semantic alignment alone is insufficient but still valuable as a supporting layer. Together with structural and legal measures, it enhances the comprehensiveness of PatentScore, ensuring that both form and meaning are adequately captured.
@@ 6 Conclusion
This research proposed PatentScore, a multi-dimensional framework that systematically integrates structural, legal, and semantic metrics for evaluating patent claims. Unlike conventional NLG metrics that fail to capture domain-specific requirements, PatentScore explicitly incorporates legal soundness and structural validity, achieving strong alignment with expert judgments ($r=0.819$) and significantly reducing error rates.
By bridging legal expertise with language modeling, PatentScore provides a reliable, transparent, and practically useful evaluation foundation, establishing a pathway for more trustworthy and innovative applications of AI in the patent domain. The experimental results show that structural and legal factors play a decisive role in assessing claim quality, while semantic metrics provide complementary value by reinforcing terminology consistency. Through expert-based validation, comparative benchmarking, and ablation studies, PatentScore is demonstrated to be a rigorous, interpretable, and reliable tool for evaluating LLM-generated claims.
While the proposed framework is developed with a focus on patent claims, it can be easily extended to other high-precision document types that require structural constraints and legal clarity, such as contract clauses, policy reports, and regulatory compliance in clinical documentation. This contributes to the broader scope of establishing domain-specific evaluation criteria for the trustworthy deployment of LLMs in high-stakes settings.
@@
This study has several limitations, which also suggest promising directions for future research.
First, the dataset used in this study is restricted to a specific technical domain and filing period. Expanding the scope to include a broader range of technologies and time frames would enhance the generalizability of the framework.
Second, PatentScore primarily evaluates intrinsic aspects of claim quality, such as structural, legal, and semantic dimensions, but does not address extrinsic properties like novelty or distinctiveness. Future work could incorporate comparative modules that measure overlap with prior art, enabling assessments of uniqueness and scope validity, which are critical in real-world patent examination.
Third, while PatentScore is designed for evaluating LLM-generated claims, its modular architecture allows adaptation to human-written patents as well. This extension would enable automated analysis and feedback on structural coherence, clarity, and legal soundness in expert-authored claims.
Fourth, integrating legal knowledge graphs into the generation and evaluation pipeline could further improve factual accuracy and contextual alignment. Such integration would strengthen the framework’s ability to handle complex dependencies across legal and technical contexts.
@@ Acknowledgments
This work was supported by the Australian Research Council (ARC) under the following projects: LP230201022 (Ethical Enterprise Representations for Personalised Sustainable Finance, Australian Research Council Linkage Grants), DP240102050 (Data Complexity and Uncertainty-Resilient Deep Variational Learning, Australian Research Council Discovery Grants), and LE240100131 (Federated Omniverse Facilities for Smart Digital Futures, Australian Research Council Linkage Infrastructure, Equipment and Facilities Grants).
We thank the creators of the HUPD dataset (CC BY License) and OpenAI for access to GPT, both of which were essential resources for this study. We are also grateful to the three evaluators for their valuable contributions, careful assessments, and voluntary participation after providing informed consent. Their expertise and feedback were instrumental in validating the evaluation framework. Finally, we acknowledge the broader research community for developing the open-source tools and prior works that served as the foundation for this study.